Solve by the elimination method
3x+4y=2
6x+8y=4
type an ordered pair

Answer by nabla(475)

If you simplify the second equation by dividing both sides by 2, you have the first equation. Hence, for all x,y that solve the first equation, the system is satisfied.
For an ordered pair, solve:
4y=2-3x
y=(1/2)-(3x/4)
Take x=1.
Then y=(1/2)-(3/4)=-1/4
So (1,-1/4) is an ordered pair.
Check it!
6(1)+8(-1/4)=6-2=4 which is the right hand side of the second equation.
Elimination means to eliminate variables, so you could just subtract two times the first equation from the second to get 0=0, but the above is a better argument.